# Emotion-Coherent Reasoning for Multimodal LLMs via Emotional Rationale Verifier

### v3

@@ 1 Introduction
Figure 1: Illustrative comparison between the baseline and our model on explanation–emotion alignment. While both correctly predict the final emotion, the baseline fails to generate an emotionally coherent explanation, highlighting its misalignment.
@@ 3.1 Model Architecture and Training Overview
Figure 2: (a): The policy model $\pi_{\theta}$ generates $G$ responses $o_{1}$, …, $o_{G}$, and the Emotional Rationale Verifier (ERV) assigns an explanation reward $R_{E}$ to each response. (b): For each response $o_{i}$, its explanation $E_{i}$ is extracted and evaluated to produce its reward $R_{i,E}$. Collectively, $R_{E}=\{R_{1,E},R_{2,E},\ldots,R_{G,E}\}$ denotes the set of explanation rewards for the $G$ responses.
@@ 3.3 Explanation Reward
Incorporating $R_{E}$ during GRPO training enhances consistency between the model’s reasoning and the target emotion.
@@ 4.2 Evaluation Metrics
Read the reasoning content and respond with the appropriate emotion in {Emotion List}.
@@ 5.1 Main Results
   W/o Neutral Filtering 46.28 42.13 64.70 57.16 38.67
W/o Sentence-Level
Verification
47.54 43.17 68.36 56.78 38.66
@@
Table 2: Ablation study of the ERV and $R_{E}$ on MAFW
@@ 5.2 Ablation Study
To validate the effectiveness of our proposed Explanation Reward, we compare it against a variant trained without our two key components: neutral filtering and sentence-level computation. In this variant, the reward calculation does not filter out non-emotional sentences (as described in Equation 4 for the neutral case), and the ERV evaluates the emotional coherence of the entire explanation holistically. We observed a performance drop when each of these components was removed individually. Specifically neutral filtering, which excludes emotionally irrelevant sentences, was particularly effective in guiding the model to generate emotion-focused explanations.
@@
Table 4: Performance of Baseline and Ours (EEA, FCR, EPC) using different closed-source LLMs.
@@ Acknowledgments
This work was partly supported by two funds: Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT) (No. RS-2025-25442384, QuestioningAI: Development of Active Reasoning Cooperative Multimodal Agent Technology), and the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No. NRF-2022R1A2C2005529).
@@ Appendix E Training Details
Number of frames
8 uniformly
sampled frames
@@
Optimizer
AdamW
($\beta_{1}=0.9$,
$\beta_{2}=0.999$)
@@
Trainable params
language model,
vision tower,
visual projector,
audio projector,
text gate module
language model,
vision tower,
visual projector,
audio projector,
text gate module
language model,
visual projector,
audio projector,
text encoder
language model,
visual projector,
audio projector,
text encoder
Vision Encoder
siglip-base
(patch16-224)
siglip-so400m
(patch14-384)
siglip-base
(patch16-224)
siglip-so400m
(patch14-384)

### v4

@@ 1 Introduction
Figure 1: Illustrative comparison between the baseline and our model on explanation–emotion alignment (text truncated for brevity). While both correctly predict the final emotion, the baseline fails to generate an emotionally coherent explanation, highlighting its misalignment.
@@ 3.1 Model Architecture and Training Overview
Figure 2: (a): The policy model $\pi_{\theta}$ generates $G$ responses $o_{1}$, …, $o_{G}$, and the Emotional Rationale Verifier (ERV) assigns an explanation reward $R_{E}$ to each response. (b): For each response $o_{i}$, its explanation $E_{i}$ is extracted and evaluated to produce its reward $R_{i,E}$. The ground truth emotion used for evaluation in this scenario is ‘Happy’($e_{gt}=$ ‘Happy’). Collectively, $R_{E}=\{R_{1,E},R_{2,E},\ldots,R_{G,E}\}$ denotes the set of explanation rewards for the $G$ responses.
@@ 3.3 Explanation Reward
Incorporating $R_{E}$ during GRPO training significantly enhances consistency between the model’s reasoning and the target emotion.
@@ 4.2 Evaluation Metrics
Read the reasoning content and respond with the appropriate emotion in
{Emotion List}.
@@ 5.1 Main Results
W/o Neutral Filt. 46.28 42.13 64.70 57.16 38.67
W/o Sent. Verif. 47.54 43.17 68.36 56.78 38.66
@@
Table 2: Ablation study of the ERV and $R_{E}$ on MAFW. All values are reported in percentage (%).
@@ 5.2 Ablation Study
To validate the effectiveness of our proposed Explanation Reward, we compare it against a variant trained without our two key components: (1) Neutral Filtering and (2) Sentence-Level Verification. In this variant, the reward calculation does not filter out non-emotional sentences (as described in Equation 4 for the neutral case), and the ERV evaluates the emotional coherence of the entire explanation holistically. We observed a performance drop when each of these components was removed individually. Specifically neutral filtering, which excludes emotionally irrelevant sentences, was particularly effective in guiding the model to generate emotion-focused explanations.
@@
Table 4: Performance of Baseline and Ours (EEA, FCR, EPC) using different closed-source LLMs. All values are reported in percentage (%).
@@ Acknowledgments
This work was partly supported by two funds: Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (No. RS-2025-25442384, QuestioningAI: Development of Active Reasoning Cooperative Multimodal Agent Technology), and the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No. NRF-2022R1A2C2005529). Additionally, this work was supported by the National Supercomputing Center with supercomputing resources including technical support (KSC-2025-CRE-0090).
@@ Appendix E Training Details
Number of frames 8 uniformly sampled frames
@@
Optimizer AdamW ($\beta_{1}=0.9$, $\beta_{2}=0.999$)
@@
Trainable params language model, vision tower, visual projector, audio projector, text gate module language model, vision tower, visual projector, audio projector, text gate module language model, visual projector, audio projector, text encoder language model, visual projector, audio projector, text encoder
Vision Encoder siglip-base (patch16-224) siglip-so400m (patch14-384) siglip-base (patch16-224) siglip-so400m (patch14-384)